The problem reads as follows:
Ideally, a school district would like the boy/girl ratio in each classroom to be essentially the same as the ratio of boys to girls in the whole class. Suppose an incoming fifth grade class is made up of 123 boys and 87 girls. Assuming there are 7 classrooms, how many boys and girls will ideally be in each room?
I am puzzled on how to complete this one, any help would be greatly appreciated.

Answer by stanbon(48545)

Ideally, a school district would like the boy/girl ratio in each classroom to be essentially the same as the ratio of boys to girls in the whole class. Suppose an incoming fifth grade class is made up of 123 boys and 87 girls. Assuming there are 7 classrooms, how many boys and girls will ideally be in each room?
------------------
Total number of 5th grade students = 123+87 = 210
The percentage of boys is 0.5857...
Assuming the 7 classes are 7 classes of 5th graders, there are 30
students in each room.
EQUATION:
Let "x" be the number of boys in each room:
x = 0.5857*30 = 17.57 (number of boys in each class)
30-x = 12.43 (number of girls in each class)
------------------
Rounding off:
18 boys in each class
12 girls in each class
================================
Cheers,
Stan H